# Parallelized Autoregressive Visual Generation

### v1

@@ Abstract
Autoregressive models have emerged as a powerful approach for visual generation but suffer from slow inference speed due to their sequential token-by-token prediction process. In this paper, we propose a simple yet effective approach for parallelized autoregressive visual generation that improves generation efficiency while preserving the advantages of autoregressive modeling. Our key insight is that parallel generation depends on visual token dependencies—tokens with weak dependencies can be generated in parallel, while strongly dependent adjacent tokens are difficult to generate together, as their independent sampling may lead to inconsistencies. Based on this observation, we develop a parallel generation strategy that generates distant tokens with weak dependencies in parallel while maintaining sequential generation for strongly dependent local tokens. Our approach can be seamlessly integrated into standard autoregressive models without modifying the architecture or tokenizer. Experiments on ImageNet and UCF-101 demonstrate that our method achieves a 3.6$\times$ speedup with comparable quality and up to 9.5$\times$ speedup with minimal quality degradation across both image and video generation tasks. We hope this work will inspire future research in efficient visual generation and unified autoregressive modeling. Project page: https://epiphqny.github.io/PAR-project.
@@ 1 Introduction
Autoregressive modeling has achieved remarkable success in language modeling [35, 36, 4, 51, 52], inspiring its application to visual generation [54, 55, 39, 33, 7, 37, 11, 65, 49, 47, 23]. These models show great potential for visual tasks due to their strong scalability and unified modeling capabilities [13, 48, 61]. Current autoregressive visual generation approaches typically rely on a sequential token-by-token generation paradigm: visual data is first encoded into token sequences using an autoencoder [56, 11, 67], then an autoregressive transformer [57] is trained to predict these tokens following a raster scan order [7]. However, this strictly sequential generation process leads to a slow generation speed, severely limiting its practical applications [27, 61]. In this work, we aim to develop an efficient autoregressive visual generation approach that improves generation speed while maintaining the generation quality.
@@ 4.1 Experimental Setup
Video Generation. We evaluate on the UCF-101 [44] dataset using our reproduced MAGVIT-v2 tokenizer [67]. Each 17-frame video (128$\times$128 resolution) is compressed by 8$\times$ spatially and 4$\times$ temporally into a $5\times 16\times 16$ token sequence (1280 tokens per video). For fair comparison, we implement both next-token prediction and our parallel generation approach using the same architecture. The position of video codes is encoded via 3D positional embeddings. Our reproduced MAGVIT-v2 tokenizer uses a 64K visual vocabulary instead of the original 262K to facilitate model training. We use Fréchet Video Distance (FVD) [53] to evaluate generation quality.
@@ 4.3 Ablation Study
Impact of attention pattern. To enable effective parallel prediction while preserving rich context modeling, we study different attention patterns between parallel predicted tokens. With $n=4$ parallel tokens, enabling full attention within groups reduces FID from 3.64 to 2.61 compared to causal attention, as it allows each token to access complete context from previous groups. This supports our design of combining bi-directional attention within groups with autoregressive attention between groups for effective parallel generation.
@@
Model scaling analysis. We study how our parallel prediction approach scales with model size. As shown in Tab. 4 (e), increasing model size from 343M to 3.1B parameters steadily improves generation quality (FID decreases from 3.76 to 2.29). Comparing with sequential generation baseline (LlamaGen) in Tab. 4.1, while smaller models show a noticeable quality gap (343M: FID 3.76 vs. 3.07), larger models achieve comparable performance (775M: 2.61 vs. 2.62; 1.4B: 2.35 vs. 2.34) while reducing generation steps from 576 to 147. This demonstrates that increased model capacity helps effectively mitigate the quality trade-off from parallel prediction, suggesting stronger capability in modeling joint distribution of parallel tokens.
@@ 5 Conclusion
We propose Parallelized Autoregressive Visual Generation (PAR), a simple yet effective approach that enables efficient parallel generation while preserving the advantages of autoregressive modeling. Our key finding is that the feasibility of parallel generation depends on token dependencies - tokens with weak dependencies can be generated in parallel while strongly dependent tokens lead to inconsistent results. Based on this insight, our PAR organizes tokens based on their dependency strengths rather than spatial proximity. The effectiveness of our approach across different visual domains validates this token dependency-based strategy for efficient autoregressive visual generation. We hope our work can inspire future research on visual generation and other sequence prediction tasks.
@@
  • Sec. B provides more visualization results.
  • Sec. C provides the analysis of visual token dependencies.
@@ A Implementation details for PAR
B More Visualization Results
In Fig.6 and Fig.7, we provide additional visualization results of PAR-4$\times$ and PAR-16$\times$ image generation on ImageNet [9] dataset, respectively.
In Fig.8, we provide the visualization results of video generation using our model on the UCF-101[44] dataset. The results are sampled from 128×128 resolution videos with 17 frames. As shown in the figure, even with 16× parallelization (PAR-16$\times$), our method shows no obvious quality degradation compared to single-token prediction (PAR-1$\times$), producing smooth motion and stable backgrounds across frames.
@@ B More Visualization Results
C Analysis of Visual Token Dependencies
In Sec.3.1, we demonstrated through pilot studies that parallel generation of adjacent tokens leads to quality degradation due to strong dependencies, while tokens from distant regions can be generated simultaneously. In this section, we provide a theoretical perspective of conditional entropy to explain this observation and our design. We use conditional entropy to measure the token dependencies quantitatively - lower conditional entropy between tokens indicates stronger dependency, while higher conditional entropy suggests weaker dependency and thus potential for parallel generation. We further validate our PAR design from the perspective of conditional entropy - In AR-based generation, each step predicts a conditional distribution of the next tokens given all previous tokens. Higher conditional entropy indicates higher difficulty for the model to predict the next tokens. In this section, we first introduce the estimation of conditional entropy in Sec.C.1, and then validate our proposed approach by analyzing the relationship between token dependencies and spatial distances in Sec. C.2.
C.1 Conditional Entropy Estimation
@@
C.2 Entropy Analysis on ImageNet Data and PAR

### v3

@@ Abstract
Autoregressive models have emerged as a powerful approach for visual generation but suffer from slow inference speed due to their sequential token-by-token prediction process. In this paper, we propose a simple yet effective approach for parallelized autoregressive visual generation that improves generation efficiency while preserving the advantages of autoregressive modeling. Our key insight is that parallel generation depends on visual token dependencies—tokens with weak dependencies can be generated in parallel, while strongly dependent adjacent tokens are difficult to generate together, as their independent sampling may lead to inconsistencies. Based on this observation, we develop a parallel generation strategy that generates distant tokens with weak dependencies in parallel while maintaining sequential generation for strongly dependent local tokens. Our approach can be seamlessly integrated into standard autoregressive models without modifying the architecture or tokenizer. Experiments on ImageNet and UCF-101 demonstrate that our method achieves a 3.6$\times$ speedup with comparable quality and up to 9.5$\times$ speedup with minimal quality degradation across both image and video generation tasks. We hope this work will inspire future research in efficient visual generation and unified autoregressive modeling. Project page: https://yuqingwang1029.github.io/PAR-project.
@@ 1 Introduction
Autoregressive modeling has achieved remarkable success in language modeling [38, 39, 4, 55, 56], inspiring its application to visual generation [58, 59, 43, 36, 8, 40, 12, 69, 53, 51, 25, 35, 7, 41]. These models show great potential for visual tasks due to their strong scalability and unified modeling capabilities [14, 52, 65]. Current autoregressive visual generation approaches typically rely on a sequential token-by-token generation paradigm: visual data is first encoded into token sequences using an autoencoder [60, 12, 71], then an autoregressive transformer [61] is trained to predict these tokens following a raster scan order [8]. However, this strictly sequential generation process leads to a slow generation speed, severely limiting its practical applications [29, 65]. In this work, we aim to develop an efficient autoregressive visual generation approach that improves generation speed while maintaining the generation quality.
@@ 4.1 Experimental Setup
Video Generation. We evaluate on the UCF-101 [48] dataset using MAGVIT-v2 tokenizer [71] reproduced by  [41]. Each 17-frame video (128$\times$128 resolution) is compressed by 8$\times$ spatially and 4$\times$ temporally into a $5\times 16\times 16$ token sequence (1280 tokens per video). For fair comparison, we implement both next-token prediction and our parallel generation approach using the same architecture. The position of video codes is encoded via 3D positional embeddings. Our reproduced MAGVIT-v2 tokenizer uses a 64K visual vocabulary instead of the original 262K to facilitate model training. We use Fréchet Video Distance (FVD) [57] to evaluate generation quality.
@@ 4.3 Ablation Study
Impact of attention pattern. To enable effective parallel prediction while preserving rich context modeling, we study different attention patterns between parallel predicted tokens. With $n=4$ parallel tokens, enabling full attention within groups reduces FID from 3.64 to 2.61 compared to causal attention, as it allows each token to access complete context from previous groups. This supports our design of combining bi-directional attention within groups with autoregressive attention between groups.
@@
Model scaling analysis. In Tab. 4(e), we study how our parallel prediction approach scales with model size. With $n=4$ parallel tokens, increasing model size from 343M to 3.1B parameters steadily improves generation quality (FID decreases from 3.76 to 2.29). Comparing with sequential generation baseline (LlamaGen) in Tab. 4.1, while smaller models show a noticeable quality gap (343M: FID 3.76 vs. 3.07), larger models achieve comparable performance (775M: 2.61 vs. 2.62; 1.4B: 2.35 vs. 2.34) while reducing generation steps from 576 to 147. This demonstrates that increased model capacity helps mitigate the quality trade-off from parallel prediction, suggesting stronger capability in modeling joint distribution of parallel tokens.
@@ 5 Conclusion
We propose PAR, a noval autoregressive visual generation approach that enables efficient parallel generation while preserving the advantages of autoregressive modeling. Our key finding is that the feasibility of parallel generation depends on token dependencies - tokens with weak dependencies can be generated in parallel while strongly dependent tokens lead to inconsistent results. Based on this insight, our PAR organizes tokens based on their dependency strengths rather than spatial proximity. The effectiveness of our approach across different visual domains validates this strategy for efficient autoregressive visual generation. We hope our work can inspire future research on visual generation and other sequence prediction tasks.
Acknowledgements
This work is partially supported by the National Nature Science Foundation of China (No. 62402406). The authors would like to thank Peize Sun, Yi Jiang, Xian Liu and Yao Teng for helpful discussions on autoregressive models.
@@ Acknowledgements
  • Sec. B demonstrates the compatibility of our approach with typical LLM engineering optimizations.
  • Sec. C provides more visualization results, including zero-shot high-resolution generation and long-range dependency examples.
  • Sec. D provides the analysis of visual token dependencies.
@@ A Implementation details for PAR
B Compatibility with Typical LLM Engineering Optimizations
We investigate whether our algorithmic parallel generation approach can complement typical engineering optimizations used in LLM inference. All experiments were conducted on a single NVIDIA A100 GPU with batch size 1, generating 384×384 resolution images. For simplicity, we only implemented PyTorch’s compile feature (leveraging CUDA graph optimization) in our PAR model. As a comparison point, we tested LlamaGen [51] with vLLM [23] optimizations, which includes both page attention and CUDA graph optimizations.
Model Resolution Optimization Latency
LlamaGen-3B 384 none 12.41s
LlamaGen-3B 384 vLLM 4.12s
PAR-3B-4x 384 none 3.46s
PAR-3B-4x 384 compile 1.15s
PAR-3B-16x 384 compile 0.43s
Table 7: Compatibility with LLM engineering optimizations. Even with just PyTorch compile optimization, our PAR approach achieves substantial speedups compared to LlamaGen with more comprehensive vLLM optimizations.
As shown in Tab. 7, our algorithmic improvements and engineering optimizations are orthogonal and provide complementary benefits. Even without any engineering optimization, PAR-3B-4x (3.46s) outperforms LlamaGen-3B with comprehensive vLLM optimizations (4.12s). When implementing just the simple CUDA graph optimization through PyTorch compile, PAR-3B-4x achieves 1.15s latency, a 3.6× improvement over optimized LlamaGen. With more aggressive parallelization, PAR-3B-16x with compile further reduces latency to 0.43s, demonstrating our approach’s flexibility in speed-quality trade-offs. These results confirm that algorithm-level optimizations (reducing sequential steps) and engineering-level optimizations (improving computational efficiency) are orthogonal approaches that, when combined, maximize generation efficiency beyond what either can achieve alone.
C More Visualization Results
Zero-shot Generation on Higher Resolutions. Fig. 6 demonstrates our model’s capability for zero-shot generation at higher resolutions (512×512) using Rotary Position Embedding [50]. Despite being trained on 384×384 images, our approach effectively maintains coherent global structures and detailed local patterns in higher resolution generation. This shows the flexibility of our parallel generation framework and its compatibility with positional encoding methods that support resolution extrapolation.
Figure 6: Zero-shot generation at 512×512 resolution. Our model successfully generates coherent high-resolution images despite being trained at 384×384 resolution.
Figure 7: Long-range dependency handling. Our method successfully maintains consistency between distant but strongly related elements (deer antlers, vehicle wheels), even when generating tokens from different spatial regions in parallel.
Long-range Dependency Handling While our approach leverages the observation that spatially distant tokens typically have weaker dependencies, certain visual elements exhibit strong long-range dependencies. Fig. 7 showcases our model’s ability to maintain consistency between distant but strongly dependent visual elements, such as symmetric features (deer antlers, vehicle wheels) and coherent structures across the image.
Addtional Image Generation Visualization. In Fig.8 and Fig.9, we provide additional visualization results of PAR-4$\times$ and PAR-16$\times$ image generation on ImageNet [10] dataset, respectively.
Addtional Video Generation Visualization. In Fig.10, we provide the visualization results of video generation using our model on the UCF-101[48] dataset. The results are sampled from 128×128 resolution videos with 17 frames. As shown in the figure, even with 16× parallelization (PAR-16$\times$), our method shows no obvious quality degradation compared to single-token prediction (PAR-1$\times$), producing smooth motion and stable backgrounds across frames.
@@ C More Visualization Results
D Analysis of Visual Token Dependencies
In Sec.3.1, we demonstrated through pilot studies that parallel generation of adjacent tokens leads to quality degradation due to strong dependencies, while tokens from distant regions can be generated simultaneously. In this section, we provide a theoretical perspective of conditional entropy to explain this observation and our design. We use conditional entropy to measure the token dependencies quantitatively - lower conditional entropy between tokens indicates stronger dependency, while higher conditional entropy suggests weaker dependency and thus potential for parallel generation. We further validate our PAR design from the perspective of conditional entropy - In AR-based generation, each step predicts a conditional distribution of the next tokens given all previous tokens. Higher conditional entropy indicates higher difficulty for the model to predict the next tokens. In this section, we first introduce the estimation of conditional entropy in Sec.D.1, and then validate our proposed approach by analyzing the relationship between token dependencies and spatial distances in Sec. D.2.
D.1 Conditional Entropy Estimation
@@
D.2 Entropy Analysis on ImageNet Data and PAR